OpenStudy (anonymous):

Finding final velocity? A canoeist on a lake throws a 0.145 kg baseball at 20.0 m/s to a friend in a different canoe. Each canoe with its contents has a mass of 120 kg, and is at rest initially. How fast is the first canoe moving after the pitch, assuming no friction with the water? I believe the answer is 0.024 m/s, can anyone confirm this?

OpenStudy (anonymous):

@Euler271 ?? I think I've seen your posts before, and you seem very helpful!

OpenStudy (anonymous):

@austinL ? Sorry if I'm not using these right. I'm new to the forum!

OpenStudy (mrdoe):

use conservation on momentum (linear)

OpenStudy (mrdoe):

of*

OpenStudy (anonymous):

What is the formula again?

OpenStudy (mrdoe):

p=mv

OpenStudy (anonymous):

20 * 0.145?

OpenStudy (mrdoe):

that will give you the momentum of the ball, yes

OpenStudy (anonymous):

And then I use a different formula, I think... M1V1 + M2V2 = Vf (m1+m2)?

OpenStudy (mrdoe):

or just use the now known momentum to calculate what velocity the boat will have

OpenStudy (mrdoe):

since the momentum from the ball will be gained by the boat when the guy catches it

OpenStudy (anonymous):

Oh, so I use the formula for velocity! I think that's what you're saying.

OpenStudy (mrdoe):

so the ball has momentum equal to \[p=0.145kg*20m/s= 2.9 N*s\] This momentum will be transfered to the boat due to the conservation on linear momentum as so \[2.9N*s = 120kg*v\] now solve for the velocity, no need to use the kinematic eq's

OpenStudy (anonymous):

41 m/s?

OpenStudy (mrdoe):

\[2.9N*s/120kg = 0.2416 m/s\]

OpenStudy (anonymous):

OH, I see!! Thank you so much!!!

OpenStudy (mrdoe):

yvw